OpenStudy (anonymous):

Show how to solve: 2cos x^2 = 1. Answers are pi/4, 3pi/4, 5pi/4, 7pi/4.

OpenStudy (anonymous):

divide by 2 as a first step to get \(\cos^2(x)=\frac{1}{2}\) then take the square root, don't forget the \(\pm\)

OpenStudy (anonymous):

get \(\cos(x)=\frac{\sqrt{2}}{2}\) or \(\cos(x)=-\frac{\sqrt{2}}{2}\) and these should look familiar

OpenStudy (shayaan_mustafa):

Satellite73 has solved it. now you find the value of x and repeats the periods then you will get your answer.\[\large x=\cos ^{-1}(\sqrt{1/2})\]\[x=\pi/4\]the cosine function is positive in the 1st and 4th quadrant. To find the second solution subtract the reference angle(2*pi) to find the solution in 4th quadrant.\[\large x=2\pi-\pi/4\]\[\large x=7\pi/4\]

OpenStudy (shayaan_mustafa):

Now again repeat the same procedure with \[\large x=\cos ^{^{-1}}(-\sqrt{1/2})\]and you will get 3pi/4 and 5pi/4 Good luck.

OpenStudy (anonymous):

@Shayaan_Mustafa: But how you derived a general solution for this? ;)

OpenStudy (shayaan_mustafa):

you mean\[\large 7\pi/4+2n \pi\]are you asking for this?

OpenStudy (anonymous):

um I got for the +ve part, \[2n\pi \pm \frac{\pi}{4} \] and for the negative part, \[2n\pi \pm \frac{3\pi}{4} \] Now how to combine this two?

OpenStudy (anonymous):

You are getting me right?

OpenStudy (shayaan_mustafa):

where are others two?

OpenStudy (anonymous):

Sorry, I don't understand.

OpenStudy (shayaan_mustafa):

and just write the full set of solution will be then x=....... so on.

OpenStudy (anonymous):

Sorry, I am not getting it :(

OpenStudy (shayaan_mustafa):

ok let me dinner then i will again online. till then kindly be patient.

OpenStudy (anonymous):

lol, really? :P

OpenStudy (anonymous):

@Shayaan_Mustafa: I was talking about the general solution not for this particular problem. And yes I got it now. Simply looking at the unit circle gives the answer.

OpenStudy (anonymous):

To be honest there is no need to do any analysis at all.

OpenStudy (anonymous):

All odd multiples of \( \frac{\pi}{4}\) gives the solution.

OpenStudy (anonymous):

Which answers he OP's query too!!